Grade 11Physical Chemistry

Copper has fcc structure with a unit cell edge of 3.61 A the size of the largest atom which could fit into interstices of the copper lattice without distorting is ?

Tanya
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
 
Copper exists in face centered cubic lattice with unit cell edge of 361pm . What is the size of the largest atom , which would fit into ocathedral voids without disturbimg the lattice
in the above the radius of Cu=a*sqrt2/4=361*1.414/4=127 pm
therrefore size of largest atom=180.5-127 pm=52.8 pm